I'm working on Adv. Alegbra II homework. I've done all of the problums exept for "Solving each Equation" questions; There are three equations that i don't understand how to solve. They are the following:
(#1) -6x^2-30=0
(#2) 3x^2+18=0
(#3) 2/3x^2+30=0
If you could help ASAP that would be awesome. :)

Answer by umnikmo(75)

All of these equations don't have real solutions, so i assume you are were introduced to i=sqrt(-1).
1. -6x^2=30
x^2=-5
X1=i*sqrt(5)
X2=-i*sqrt(5)
2. 3x^2=-18
x1=i*sqrt(6)
x2=-i*sqrt(6)
3. x^2=-45
x1=3i*sqrt(5)
x2=-3i*sqrt(5)